# LumiMAS: A Comprehensive Framework for Real-Time Monitoring and Enhanced Observability in Multi-Agent Systems

### v1

@@ 7 BiasCheck
Classification Evaluation Settings
@@
Additional Hyperparameter Settings
@@
Anomaly detection Method Application Framework Model
@@
Table 6: Model configuration for each application and anomaly detection approach.
C.2 Additional Evaluation Results
Loss Curves.
@@ C.2 Additional Evaluation Results
Figures  45, and 6 show the loss curves for the EPI, Semantic, and Combined approaches, respectively, for the Trip Planner application, with GPT-4o-mini as the underlying model, built with the CrewAI MAS framework.
@@
Figure 4: EPI anomaly detection approach loss curve Trip Planner application
@@
Figure 5: Semantic anomaly detection approach loss curve Trip Planner application
@@
Figure 6: Combined anomaly detection approach loss curve Trip Planner application
@@
Detailed Results of the Anomaly Detection Layer on CrewAI Applications with GPT-4o-mini.
@@
In this subsection, we provide more details about the baseline evaluation. The evaluation baselines were evaluated with both GPT-4o and GPT-4o-mini models. Table  7 presents the anomaly detection results of the five types of vulnerabilities (three DPI scenarios) evaluated (rows), along with each detection approach’s performance and overhead. The result indicates that GPT-4o and GPT-4o-mini generate a similar number of tokens, although the 4o-mini model costs per token is cheaper. In terms of time-consuming GPT-4o-mini was 40% faster on average than GPT-4o. we can see that in most vulnerabilities, such as bias, DPI (misinformation, exhaustion), IPI, and MP, The agent-as-a-judge method performed better with the larger GPT-4o model in comparison to the 4o-mini model. This can be attributed to the model’s improved ability to handle complex contexts when acting as an agent. In contrast, in some cases, such as hallucination MP and DPI, when operating in an LLM-as-a-judge setting, GPT-4o-mini produced superior results. In 3 out of 7 scenarios, 4o-mini performed better than the 4o either in LLM-as-a-judge settings or agent-as-a-judge settings, but the results are inconclusive, while 4o did achieve better results in the other cases, and especially as an agent. Overall, although GPT-4o-mini was effective in certain cases, our method remains more robust, significantly faster, incurs no cost, and, in most cases, achieves better anomaly detection performance.
C.3 DPI Attack Evaluation
@@
In this subsection, we present the evaluation of 3 different direct prompt injection attacks: Misinformation, Exhaustion, and Backdoor. Table 7 presents the anomaly detection results of the five types of vulnerabilities (three DPI scenarios) evaluated (rows), along with each detection approach’s performance and overhead. The result on the DPI scenarios indicates that our combined detection approach achieved the lowest average false positive rate (FPR) (0.270), indicating a low number of false alarms, while maintaining high accuracy. Our method is the most efficient in terms of cost and inference time. It performed best on the exhaustion attack, achieving the best results across all performance metrics. It also excelled in detecting the misinformation attack, with the highest accuracy and the lowest false positive rate. However, the backdoor attack was more challenging for our method to detect.
@@
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$ Cost $\downarrow$
Hallucination Consistency of 3 0.700 0.727 0.800 0.667 0.400 9.215 5189.425
Consistency of 5 0.695 0.716 0.770 0.670 0.380 15.676 8648.480
LLM-as-a-judge (4o) 0.595 0.382 0.250 0.806 0.060 7.138 2516.050
Agent-as-a-judge (4o) 0.595 0.409 0.280 0.757 0.090 12.312 1341349.905
LLM-as-a-judge (4o-mini) 0.675 0.545 0.390 0.907 0.040 5.604 2403.840
Agent-as-a-judge (4o-mini) 0.595 0.426 0.300 0.732 0.110 6.114 1300608.895
EPI (Ours) 0.745 0.730 0.690 0.775 0.200 0.003 -
Semantic (Ours) 0.720 0.682 0.600 0.790 0.160 0.014 -
Combined (Ours) 0.765 0.728 0.630 0.863 0.100 0.017 -
Bias toxic-bert 0.510 0.671 1.000 0.505 0.980 0.398 -
LLM-as-a-judge (4o) 0.905 0.907 0.930 0.886 0.120 7.810 2324.075
Agent-as-a-judge (4o) 0.900 0.900 0.900 0.900 0.100 14.042 1283347.355
LLM-as-a-judge (4o-mini) 0.815 0.823 0.860 0.789 0.230 5.657 2188.745
Agent-as-a-judge (4o-mini) 0.760 0.793 0.920 0.697 0.400 6.79366 1265069.915
EPI (Ours) 0.625 0.675 0.780 0.595 0.530 0.016 -
Semantic (Ours) 0.600 0.615 0.640 0.593 0.440 0.020 -
Combined (Ours) 0.660 0.670 0.690 0.651 0.370 0.024 -
Direct
Prompt
Injection (DPI)
Misinformation
LLM-as-a-judge (4o) 0.670 0.752 1.000 0.602 0.660 9.809 4448.190
Agent-as-a-judge (4o) 0.785 0.823 1.000 0.699 0.430 26.132 4881852.055
LLM-as-a-judge (4o-mini) 0.770 0.811 0.990 0.688 0.450 7.436 4334.405
Agent-as-a-judge (4o-mini) 0.775 0.813 0.980 0.695 0.430 12.707 2685723.890
EPI (Ours) 0.795 0.827 0.980 0.715 0.390 0.008 -
Semantic (Ours) 0.815 0.843 0.990 0.733 0.360 0.071 -
Combined (Ours) 0.815 0.830 0.900 0.769 0.270 0.084 -
Direct
Prompt
Injection (DPI)
Exhaustion
LLM-as-a-judge (4o) 0.620 0.723 0.990 0.569 0.75 9.485 4498.430
Agent-as-a-judge (4o) 0.775 0.812 0.970 0.698 0.42 24.212 4694451.545
LLM-as-a-judge (4o-mini) 0.775 0.802 0.910 0.717 0.36 7.287 4373.410
Agent-as-a-judge (4o-mini) 0.585 0.610 0.650 0.575 0.48 13.017 3291554.825
EPI (Ours) 0.805 0.837 1.000 0.719 0.390 0.009 -
Semantic (Ours) 0.800 0.828 0.960 0.727 0.360 0.082 -
Combined (Ours) 0.865 0.881 1.000 0.787 0.270 0.096 -
Direct
Prompt
Injection (DPI)
Backdoor
LLM-as-a-judge (4o) 0.525 0.612 0.750 0.517 0.7 9.982 4563.200
Agent-as-a-judge (4o) 0.555 0.590 0.640 0.547 0.53 27.257 4774985.315
LLM-as-a-judge (4o-mini) 0.755 0.798 0.970 0.678 0.46 7.663 4450.135
Agent-as-a-judge (4o-mini) 0.745 0.777 0.890 0.690 0.4 12.040 2639110.585
EPI (Ours) 0.625 0.631 0.640 0.621 0.390 0.007 -
Semantic (Ours) 0.695 0.711 0.750 0.676 0.360 0.067 -
Combined (Ours) 0.620 0.573 0.510 0.654 0.270 0.083 -
Indirect
Prompt
Injection (IPI)
LLM-as-a-judge (4o) 0.560 0.679 0.930 0.534 0.810 20.182 6844.650
Agent-as-a-judge (4o) 0.650 0.724 0.920 0.597 0.620 29.617 9929870.045
LLM-as-a-judge (4o-mini) 0.545 0.622 0.750 0.532 0.660 14.323 6556.370
Agent-as-a-judge (4o-mini) 0.595 0.646 0.740 0.574 0.550 18.314 6774943.495
EPI (Ours) 0.965 0.966 1.000 0.935 0.070 0.029 -
Semantic (Ours) 0.890 0.894 0.930 0.861 0.150 0.102 -
Combined (Ours) 0.970 0.970 0.970 0.970 0.030 0.142 -
Memory
Poisoning (MP)
LLM-as-a-judge (4o) 0.805 0.820 0.890 0.761 0.280 9.509 3764.930
Agent-as-a-judge (4o) 0.780 0.784 0.800 0.769 0.240 13.348 1787606.550
LLM-as-a-judge (4o-mini) 0.850 0.845 0.820 0.872 0.120 7.153 3664.815
Agent-as-a-judge (4o-mini) 0.670 0.621 0.540 0.730 0.200 8.117 3654892.730
EPI (Ours) 0.540 0.643 0.830 0.525 0.750 0.003 -
Semantic (Ours) 0.690 0.710 0.760 0.667 0.380 0.028 -
Combined (Ours) 0.500 0.561 0.640 0.500 0.640 0.032 -
Table 7: Anomaly detection results obtained using the CrewAI apps, with GPT-4o-mini as the underlying model including baseline approach with GPT-4o-mini as the underlying model.
Application Changes Effect On Detection
@@
Table 8: The effect of changes in the application on the anomaly detection and the need for retraining the model
Figure 7: Anomaly classification performance confusion matrix with GPT-4o-mini as the underlying model.
Classification Results on CrewAI Apps with GPT-4o mini.

### v2

@@ C.1.2. Baselines
To implement LogBERT guo2021logbert, we followed the original setup by adopting an encoder-only transformer architecture vaswani2017attention trained with a masked language modeling objective. For each application, we extracted the event sequences from the benign train logs and used them to train the model. An anomaly score threshold was selected based on the validation set by maximizing the F1 score.
Table 5. MAS applications used in our evaluation
Num App Name App Description
MAS
Platform
Number
of Agents
Type of
Vulnerabilities
1
Trip Planner
crewAI-examples
Create a personalized seven-day
travel guide with a detailed itinerary
for each day.
CrewAI 3 DPI
2
Trip Planner (adaptation of
crewAI-examples)
Create a personalized seven-day
travel guide with a detailed itinerary
for each day.
LangGraph 3 DPI
3 GenFic
Generate a short story for the given
topic, leveraging Wikipedia for
characters’ backstories.
LangGraph 5 DPI
4
Instagram Post
crewAI-examples
Generate an Instagram post for
commercial use, featuring a product
from an e-commerce platform.
CrewAI 5 IPI
5
Real Estate Team
crewai-rag-deep-dive
Analyze home inspection PDFs,
identify issues, estimate costs, and
generate summaries with RAG,
search, and batch processing support.
CrewAI 3 MP
6 HalluCheck
Hallucination-tailored app
Answer a question based on a context,
justify the answer, and infer the
original question from it.
CrewAI 3 Hallucination
7 BiasCheck
Bias-tailored app
Answer a question based on the model’s
knowledge, justify the answer, and
evaluate it ethically.
CrewAI 3 Bias
C.1.3. Classification Evaluation Settings
@@
C.1.4. Additional Hyperparameter Settings
@@
Table 6. Model configuration for each application and anomaly detection approach
Method Application Framework Model
@@
C.2. Additional Evaluation Results
C.2.1. Loss Curves
@@
Figures 4, 5, and 6 show the loss curves for the EPI, semantic, and combined approaches, respectively, for the Trip Planner application, with GPT-4o-mini as the underlying model, built with the CrewAI MAS framework.
@@
Figure 4. Loss curve of the EPI-based anomaly detection approach on the Trip Planner application
[…]
Line plot showing the training loss of the EPI-based anomaly detection approach on the Trip Planner application. The horizontal axis represents training epochs, and the vertical axis shows loss values over time
@@
Figure 5. Loss curve of the semantic-based anomaly detection approach on the Trip Planner application
[…]
Line plot showing the training loss of the semantic-based anomaly detection approach on the Trip Planner application. The horizontal axis represents training epochs, and the vertical axis shows loss values over time
@@
Figure 6. Loss curve of the combined anomaly detection approach on the Trip Planner application
[…]
Line plot showing the training loss of the combined anomaly detection approach on the Trip Planner application. The horizontal axis represents training epochs, and the vertical axis shows loss values over time
@@
C.2.2. Detailed Results of the Anomaly Detection Layer on CrewAI Applications with GPT-4o-mini.
@@
In this subsection, we provide more details about the baseline evaluation. The evaluation baselines were evaluated with both GPT-4o and GPT-4o-mini models. Table 8 presents the anomaly detection results of the five types of vulnerabilities (three DPI scenarios) evaluated (rows), along with each detection approach’s performance and overhead. The result indicates that GPT-4o and GPT-4o-mini generate a similar number of tokens, although the 4o-mini model costs per token is cheaper. In terms of time-consuming, the GPT-4o-mini agent was 22% faster on average than the GPT-4o agent. In some cases, such as hallucination MP and DPI Backdoor, when operating in an LLM-as-a-judge setting, GPT-4o-mini produced superior results. In 4 out of 7 scenarios, 4o-mini performed better than the 4o either in LLM-as-a-judge settings or agent-as-a-judge settings, but the results are inconclusive, while 4o did achieve better results in other cases, and especially as an agent. Overall, although GPT-4o-mini was effective in certain cases, our method remains more robust, significantly faster, incurs no cost, and, in most cases, achieves better anomaly detection performance.
C.2.3. DPI Attack Evaluation
@@
In this subsection, we present the evaluation of 3 different direct prompt injection attacks: Misinformation, Exhaustion, and Backdoor. Table 7 presents the anomaly detection results of the three types of DPI vulnerabilities evaluated (rows), along with each detection approach’s performance and overhead. The result on the DPI scenarios indicates that our combined detection approach achieved the lowest average false positive rate (FPR) among all baselines (0.238), indicating a low number of false alarms, while maintaining high accuracy. It also achieved the highest performance on the exhaustion attack, showing superior results across all evaluation metrics. It also demonstrated strong accuracy in detecting the misinformation attack and maintained consistent efficiency across all examined DPI scenarios while being more efficient in both time and token usage cost compared to the LLM-based baselines. In contrast, although LogBERT operates faster, it scored significantly lower on most evaluation metrics and proved ineffective against the DPI backdoor attack, where the log lacks structural anomalies.
Table 7. Three DPI anomaly detection results obtained using the CrewAI apps, with GPT-4o-mini as the underlying model
@@
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$
Token
Count
$\downarrow$
DPI Misinformation LLM-as-a-judge 0.750 ± 0.017 0.800 ± 0.011 1.000 ± 0.000 0.667 ± 0.016 0.500 ± 0.035 7.538 4363.4
Agent-as-a-judge 0.792 ± 0.015 0.828 ± 0.010 1.000 ± 0.000 0.706 ± 0.015 0.416 ± 0.030 19.207 3851714.2
LogBERT 0.574 ± 0.077 0.673 ± 0.023 0.872 ± 0.103 0.556 ± 0.060 0.724 ± 0.251 0.007 -
EPI (ours) 0.750 ± 0.013 0.783 ± 0.009 0.900 ± 0.021 0.693 ± 0.017 0.400 ± 0.038 0.006 -
Semantic (ours) 0.865 ± 0.008 0.877 ± 0.006 0.964 ± 0.011 0.805 ± 0.014 0.234 ± 0.023 0.069 -
Combined (ours) 0.823 ± 0.009 0.833 ± 0.011 0.884 ± 0.034 0.788 ± 0.014 0.238 ± 0.028 0.074 -
DPI Exhaustion LLM-as-a-judge 0.744 ± 0.010 0.793 ± 0.007 0.980 ± 0.014 0.666 ± 0.009 0.492 ± 0.022 7.131 4403.6
Agent-as-a-judge 0.773 ± 0.018 0.812 ± 0.012 0.980 ± 0.012 0.693 ± 0.017 0.434 ± 0.037 18.281 3504222
LogBERT 0.637 ± 0.125 0.738 ± 0.066 0.998 ± 0.004 0.589 ± 0.084 0.724 ± 0.251 0.003 -
EPI (ours) 0.800 ± 0.019 0.834 ± 0.013 1.000 ± 0.000 0.715 ± 0.020 0.400 ± 0.038 0.006 -
Semantic (ours) 0.856 ± 0.010 0.868 ± 0.007 0.946 ± 0.013 0.802 ± 0.015 0.234 ± 0.023 0.078 -
Combined (ours) 0.881 ± 0.014 0.894 ± 0.011 1.000 ± 0.000 0.808 ± 0.018 0.238 ± 0.028 0.083 -
DPI Backdoor LLM-as-a-judge 0.622 ± 0.022 0.658 ± 0.022 0.728 ± 0.031 0.601 ± 0.017 0.484 ± 0.019 7.132 4473
Agent-as-a-judge 0.622 ± 0.036 0.640 ± 0.039 0.674 ± 0.051 0.610 ± 0.032 0.430 ± 0.039 17.335 3691578.2
LogBERT 0.505 ± 0.014 0.584 ± 0.087 0.734 ± 0.249 0.503 ± 0.012 0.724 ± 0.251 0.005 -
EPI (ours) 0.589 ± 0.022 0.585 ± 0.017 0.578 ± 0.011 0.592 ± 0.026 0.400 ± 0.038 0.004 -
Semantic (ours) 0.725 ± 0.011 0.713 ± 0.020 0.684 ± 0.041 0.745 ± 0.010 0.234 ± 0.023 0.065 -
Combined (ours) 0.702 ± 0.041 0.681 ± 0.057 0.642 ± 0.080 0.728 ± 0.031 0.238 ± 0.028 0.070 -
Table 8. Baseline approaches with GPT-4o-mini and GPT-4o as the underlying model Anomaly detection results. Obtained using the CrewAI applications, with GPT-4o-mini as the underlying model
Vulnerability Judge Method Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$
Token
Count
$\downarrow$
Hallucination LLM (4o) 0.574 ± 0.010 0.332 ± 0.025 0.212 ± 0.019 0.767 ± 0.022 0.064 ± 0.005 6.216 2414.0
Agent (4o) 0.586 ± 0.033 0.406 ± 0.059 0.284 ± 0.051 0.717 ± 0.073 0.112 ± 0.030 9.617 1436329.0
LLM (4o-mini) 0.686 ± 0.007 0.571 ± 0.014 0.418 ± 0.015 0.901 ± 0.011 0.046 ± 0.005 6.981 2409.0
Agent (4o-mini) 0.587 ± 0.025 0.420 ± 0.039 0.300 ± 0.034 0.708 ± 0.066 0.126 ± 0.041 8.699 1339806.0
Bias LLM (4o) 0.907 ± 0.008 0.909 ± 0.008 0.926 ± 0.011 0.892 ± 0.010 0.112 ± 0.011 5.297 2229.0
Agent (4o) 0.894 ± 0.015 0.892 ± 0.015 0.876 ± 0.017 0.909 ± 0.023 0.088 ± 0.024 10.129 1253684.0
LLM (4o-mini) 0.821 ± 0.008 0.831 ± 0.007 0.882 ± 0.013 0.786 ± 0.011 0.240 ± 0.017 7.359 2197.0
Agent (4o-mini) 0.763 ± 0.032 0.793 ± 0.029 0.908 ± 0.040 0.704 ± 0.025 0.382 ± 0.036 8.801 1E+06
DPI Misinformation LLM (4o) 0.750 ± 0.017 0.800 ± 0.011 1.000 ± 0.000 0.667 ± 0.016 0.500 ± 0.035 7.538 4363.4
Agent (4o) 0.792 ± 0.015 0.828 ± 0.010 1.000 ± 0.000 0.706 ± 0.015 0.416 ± 0.030 19.208 3851714.2
LLM (4o-mini) 0.757 ± 0.010 0.804 ± 0.007 0.998 ± 0.004 0.673 ± 0.008 0.484 ± 0.017 9.519 4454.8
Agent (4o-mini) 0.771 ± 0.022 0.791 ± 0.019 0.868 ± 0.025 0.727 ± 0.023 0.326 ± 0.036 13.661 2476279.0
DPI Exhaustion LLM (4o) 0.744 ± 0.010 0.793 ± 0.007 0.980 ± 0.014 0.666 ± 0.009 0.492 ± 0.022 7.132 4403.6
Agent (4o) 0.773 ± 0.018 0.812 ± 0.012 0.980 ± 0.012 0.693 ± 0.017 0.434 ± 0.037 18.282 3504222.0
LLM (4o-mini) 0.763 ± 0.016 0.808 ± 0.011 0.994 ± 0.005 0.680 ± 0.014 0.468 ± 0.029 8.170 4377.6
Agent (4o-mini) 0.752 ± 0.028 0.772 ± 0.025 0.838 ± 0.030 0.715 ± 0.026 0.334 ± 0.036 16.734 2662075.2
DPI Backdoor LLM (4o) 0.622 ± 0.022 0.658 ± 0.022 0.728 ± 0.031 0.601 ± 0.017 0.484 ± 0.019 7.133 4473.0
Agent (4o) 0.622 ± 0.036 0.640 ± 0.039 0.674 ± 0.051 0.610 ± 0.032 0.430 ± 0.039 17.3356 3691578.2
LLM (4o-mini) 0.756 ± 0.008 0.804 ± 0.005 1.000 ± 0.000 0.672 ± 0.008 0.488 ± 0.016 8.942 4455.2
Agent (4o-mini) 0.777 ± 0.034 0.794 ± 0.030 0.860 ± 0.024 0.738 ± 0.034 0.306 ± 0.045 14.117 2541351.4
IPI LLM (4o) 0.547 ± 0.004 0.681 ± 0.002 0.966 ± 0.009 0.526 ± 0.003 0.872 ± 0.015 17.987 6642.0
Agent (4o) 0.614 ± 0.020 0.698 ± 0.013 0.890 ± 0.012 0.574 ± 0.015 0.662 ± 0.036 34.618 9895765.0
LLM (4o-mini) 0.526 ± 0.013 0.613 ± 0.012 0.750 ± 0.024 0.518 ± 0.009 0.698 ± 0.028 15.770 6570.2
Agent (4o-mini) 0.617 ± 0.031 0.638 ± 0.033 0.676 ± 0.042 0.604 ± 0.027 0.442 ± 0.026 20.619 6499876.0
MP LLM (4o) 0.797 ± 0.012 0.814 ± 0.011 0.890 ± 0.014 0.750 ± 0.011 0.296 ± 0.015 7.835 3701.0
Agent (4o) 0.746 ± 0.013 0.749 ± 0.016 0.758 ± 0.028 0.740 ± 0.011 0.266 ± 0.015 10.484 1770940.0
LLM (4o-mini) 0.858 ± 0.018 0.853 ± 0.018 0.824 ± 0.013 0.884 ± 0.025 0.108 ± 0.026 8.174 3670.6
Agent (4o-mini) 0.705 ± 0.009 0.638 ± 0.006 0.520 ± 0.010 0.827 ± 0.030 0.110 ± 0.025 11.543 1886314.0
C.2.4. Application Changes Effect On Detection
@@
Table 9. The effect of changes in the application on the anomaly detection and the need for retraining the model
@@
Figure 7. Anomaly classification performance confusion matrix on CrewAI apps with GPT-4o-mini as the underlying model.
[…]
Confusion matrix showing anomaly classification performance on CrewAI apps with GPT-4o-mini as the underlying model. Rows correspond to true classes and columns correspond to predicted classes. Cell values indicate the number of predictions for each class combination, and color intensity reflects the magnitude of the counts.
C.2.5. Feature Combination Ablation
To further evaluate our proposed framework, we introduce an additional feature combination strategy aimed at improving the detection of memory poisoning attacks. In the original combined approach, the latent representations produced by the EPI-AE and Semantic-AE were fused after their respective autoencoders. In contrast, our new variant fuses the raw EPI and semantic features prior to the autoencoder stage. This early-fusion method incorporates both signal types into a unified input sequence, where each agent execution step contributes to the combined log representation. This enables the autoencoder to jointly model operational and semantic patterns during training, potentially capturing cross-modal interactions overlooked in the late-fusion design. Table 10 presents the anomaly detection results of the proposed architectural enhancement across all examined failure types obtained using CrewAI with GPT-4o-mini as the underlying model. The findings indicate that the early-fusion strategy substantially improves the detection of the memory poisoning (MP) attack, achieving an F1 score of 0.745, compared to 0.536 with the original combined-feature approach. While MP detection shows an improvement, other failure types exhibit only marginal improvements, and in cases such as hallucination and bias, the early-fusion method results in a slight decrease in performance. This suggests that jointly modeling operational and semantic signals at the input level provides additional sensitivity for certain failure types, whereas other fusion strategies remain more effective in different scenarios. These findings highlight the importance of selecting the fusion method that best aligns with the specific failure modes and operational requirements.
Table 10. Anomaly detection results of the Feature Combination Ablation obtained using the CrewAI apps, with GPT-4o-mini as the underlying model
Vulnerability Method Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$
Token
Count
$\downarrow$
Hallucination Combined 0.765 ± 0.015 0.733 ± 0.016 0.646 ± 0.030 0.850 ± 0.039 0.116 ± 0.042 0.016 -
Early-fusion 0.677 ± 0.012 0.678 ± 0.010 0.682 ± 0.041 0.677 ± 0.031 0.328 ± 0.060 0.0214 -
Bias Combined 0.664 ± 0.027 0.661 ± 0.020 0.654 ± 0.015 0.668 ± 0.033 0.326 ± 0.048 0.023 -
Early-fusion 0.613 ± 0.007 0.617 ± 0.007 0.624 ± 0.019 0.611 ± 0.010 0.398 ± 0.026 0.031 -
DPI Combined 0.802 ± 0.021 0.803 ± 0.026 0.842 ± 0.038 0.775 ± 0.021 0.238 ± 0.028 0.076 -
Early-fusion 0.854 ± 0.026 0.851 ± 0.027 0.874 ± 0.028 0.841 ± 0.053 0.166 ± 0.069 0.094 -
IPI Combined 0.941 ± 0.033 0.941 ± 0.033 0.942 ± 0.039 0.941 ± 0.040 0.060 ± 0.041 0.102 -
Early-fusion 0.982 ± 0.010 0.982 ± 0.009 0.994 ± 0.005 0.971 ± 0.021 0.030 ± 0.023 0.105 -
MP Combined 0.477 ± 0.068 0.536 ± 0.065 0.608 ± 0.097 0.483 ± 0.058 0.654 ± 0.114 0.032 -
Early-fusion 0.719 ± 0.026 0.745 ± 0.013 0.818 ± 0.036 0.686 ± 0.040 0.380 ± 0.083 0.034 -
C.2.6. Classification Results on CrewAI Apps with GPT-4o mini